OpenStudy (anonymous):

Find a quadratic equation whose roots are the reciprocals of the roots of 3x^2+5x-2=0.

OpenStudy (anonymous):

A + B = -5/3 where A and B are the roots of this equation AB = 2/3 1/A + 1/B = (A + B) / AB = -5/3 / 2/3 = -5/2 thats part of now we need the va;ue of 1/A x 1/B

OpenStudy (anonymous):

Roots of the given equation are -2 and 1/3 3x^2+5x-2=0. 3x^2 + 6x -x -2 =0 3x(x+2) - 1(x+2) = 0 3x-1) (x+2) = 0 x = 1/3 or -2 Now the reciprocals are 3 and -1/2 To find the equation we use the formula x^2 - S x + P = 0 Where S is the sum of roots ( = 3 + (-1/2)) and P is product of roots which is -3/2 Get your answer!

OpenStudy (anonymous):

this equals 1 / AB = 3/2 equation is 2x^2 - 5x + 3 = 0

OpenStudy (anonymous):

which is right?

OpenStudy (anonymous):

the rules i used are for ax^2 + bx + c = 0 if A and B are the roots then A + B = -b/a and AB = c/a i havent checked them yet - i might have made a typo of course.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yes i have made a mistake sorry AB should be -2/3 check second line of my first post - sorry

OpenStudy (anonymous):

the method is correct though

OpenStudy (anonymous):

is Lfcmaniac correct?

OpenStudy (anonymous):

i'm checking now

OpenStudy (anonymous):

yes the answer is 2x^2 - 5x -3 = 0 whose roots are 3 and -0.5

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

we got there in the end!

OpenStudy (anonymous):

how did he get this 3x^2 + 6x -x -2 =0

OpenStudy (anonymous):

jimmyrep

OpenStudy (anonymous):

ok - he multiplied the 3x (from 3x^2) by 2 (from -2) to get the 6x.Now the middle term is +5x so he took wrote - x (to get the +5x). he was aiming for +5x. If the middle term was -5x, he would have written -6x + x.

OpenStudy (anonymous):

if the quadratic will factor his was the easiest way of solving this problem. My way was more general and can be used for quadratics that will not factor.

OpenStudy (anonymous):

@jimmyrep, your method should always work, i just proved it from the quadratic formula.

OpenStudy (anonymous):

mine was not the only quadratic that fits the bill in this problem -5x^2 + 5x + 3 = 0 would also be an answer so would 10x^2 - 10x - 6 = 0 but the question asked for 'a quadratic' so there are many answers

OpenStudy (anonymous):

ty mike0